OpenStudy (anonymous):

what is the number of solutions to the system? someone please helppp mee!! please! x+ y = -1 5x- 5y = 7

OpenStudy (anonymous):

x = 1/5, y = -6/5

OpenStudy (anonymous):

so theres one solution

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

You might have multiple solutions if you were taking the square root of something but not this time.

OpenStudy (anonymous):

If you want to solve problems faster during the SAT test math sections, try our SAT app for your calculator: http://download.cnet.com/SAT-Operating-System-For-TI-83-And-TI-84/3000-20415_4-75448635.html?part=dl-&subj=dl&tag=button

OpenStudy (anonymous):

thanks but im only in 8th grade. i have to do high school math. to prepare my self for freshman year.

OpenStudy (anonymous):

No problem, but you can use it to check homework.

OpenStudy (anonymous):

ahhh okayy. thank you so much!

OpenStudy (anonymous):

No problem. Don't forget to medal! :)